### v1

# Bridging Constraints and Stochasticity: A Fully First-Order Method for Stochastic Bilevel Optimization with Linear Constraints

Abstract
This work provides the first finite-time convergence guarantees for linearly constrained stochastic bilevel optimization using only first-order methods, requiring solely gradient information without any Hessian computations or second-order derivatives. We address the unprecedented challenge of simultaneously handling linear constraints, stochastic noise, and finite-time analysis in bilevel optimization, a combination that has remained theoretically intractable until now. While existing approaches either require second-order information, handle only unconstrained stochastic problems, or provide merely asymptotic convergence results, our method achieves finite-time guarantees using gradient-based techniques alone. We develop a novel framework that constructs hypergradient approximations via smoothed penalty functions, using approximate primal and dual solutions to overcome the fundamental challenges posed by the interaction between linear constraints and stochastic noise. Our theoretical analysis provides explicit finite-time bounds on the bias and variance of the hypergradient estimator, demonstrating how approximation errors interact with stochastic perturbations. We prove that our first-order algorithm converges to $(\delta,\epsilon)$-Goldstein stationary points using $\Theta(\delta^{-1}\epsilon^{-5})$ stochastic gradient evaluations, establishing the first finite-time complexity result for this challenging problem class and representing a significant theoretical breakthrough in constrained stochastic bilevel optimization.
1 Introduction
Bilevel optimization is a powerful paradigm for hierarchical decision-making in machine learning, including hyperparameter tuning franceschi2018, meta-learning finn2017model, and reinforcement learning konda1999actor.
The standard formulation can be written as the following optimization problem:
$\min_{x\in X}f(x,y^{*}(x))\quad\text{s.t.}~y^{*}(x)\in\operatorname*{arg\,min}_{y\in S(x)}g(x,y),$ (1)
where we define $F(x)\coloneqq f(x,y^{*}(x))$ to be the overall bilevel objective as a function of $x$. Traditional methods of solving bilevel optimization often face scalability challenges pedregosa2016hyperparameter; franceschi2017forward, including implicit differentiation with its requisite Hessian computations amos2017optnet; ji2021bilevel; khanduri2024doubly; hu2023contextual; ghadimi2018approximation, and iterative differentiation characterized by high memory and computational demands. shen2024memory; brauer2024learning
A recent breakthrough in bilevel optimization by kwon2023; liu2022bome proposes using reformulation and penalty-based approaches to design a fully first-order gradient proxy. Several follow-up works based on this breakthrough have emerged, including improving finite time convergence of unconstrained bilevel optimization chen2024finding; yang2023achieving; chen2024optimal; kwon2024complexity, constrained bilevel optimization khanduri2023linearly; kornowski2024; yao2024constrained; lu2024first, and applications of bilevel algorithms to machine learning pan2024scalebio; zhang2024introduction; petrulionyte2024functional.
However, many real-world scenarios involve stochastic objectives and constraints together, where gradients are noisy estimates from samples. While methods for unconstrained stochastic bilevel optimization have advanced (e.g., ghadimi2018approximation; kwon2023; liu2022bome), the confluence of stochasticity and explicit LL linear constraints poses significant unresolved challenges. A critical gap persists: no existing methods offer finite-time convergence guarantees for bilevel problems that are simultaneously stochastic and linearly constrained in the lower-level problem.
This paper bridges this gap by introducing the Fully First-order Constrained Stochastic Approximation (F2CSA) algorithm. We build upon the deterministic framework of kornowski2024, developing a novel smoothed, stochastic hypergradient oracle tailored for bilevel problems with linearly constrained LL subproblems and stochastic objectives. The key to our approach is a smoothed reformulation that handles inexact dual variables, enabling robust hypergradient estimation from noisy first-order information and inexact LL solves. Our theoretical analysis, based on Lipschitz continuity and careful variance-bias tradeoffs, yields the first finite-time complexity guarantees for reaching $(\delta,\epsilon)$-Goldstein stationary points in this setting.
Our contributions include:
  • Stochastic Inexact Hypergradient Oracle: We develop a stochastic inexact hypergradient oracle based on a smoothed Lagrangian method with penalty weights $\alpha_{1}=\alpha^{-2}$ and $\alpha_{2}=\alpha^{-4}$ where $\alpha>0$. This oracle approximates hypergradients with bias bounded by $O(\alpha)$ and variance bounded by $O(1/N_{g})$ using $N_{g}$ first-order gradient evaluations. Our smoothed Lagrangian method generalizes the approach from kornowski2024 to allow approximate primal-dual lower-level solutions for constructing inexact hypergradient oracles.
  • Convergence Guarantees: We apply the stochastic inexact hypergradient oracle with parameter $\alpha=\epsilon$ and $N_{g}=O(\epsilon^{-2})$ to design a double-loop algorithm for stochastic bilevel optimization problems with linear constraints. This algorithm attains a $(\delta,\epsilon)$-Goldstein stationary point of $F(x)$ with total first-order oracle complexity $\tilde{O}(\delta^{-1}\epsilon^{-5})$. This generalizes the deterministic bilevel optimization result with linear constraints (rate $\tilde{O}(\delta^{-1}\epsilon^{-4})$) to the stochastic setting.
Our work provides the first finite-time convergence guarantees for linearly constrained stochastic bilevel optimization under standard stochastic assumptions, providing a theoretically sound yet practical alternative to traditional bilevel optimization approaches.
2 Related Work
Penalty Methods and First-Order Reformulations. To reduce the computational cost of second-order derivatives in bilevel optimization, recent works have proposed scalable, first-order algorithms based on penalty formulations kwon2023; liu2022bome. These techniques transform constrained bilevel problems into single-level surrogates that can be solved efficiently with convergence guarantees, where deterministic, partially stochastic, and fully stochastic bilevel optimization can achieve $\epsilon$-stationary point in $O(\epsilon^{-2})$, $O(\epsilon^{-4})$, $O(\epsilon^{-6})$ gradient calls, respectively chen2024optimal. The convergence rate of the deterministic case can be further improved to $O(\epsilon^{-1.5})$ by momentum-based method yang2023achieving.
Bilevel Optimization with Linear Constraints.
Due to the nonsmoothness of constrained bilevel optimization problems,  kornowski2024 focuses on Goldstein stationary goldstein1977 and designs a new penalty method to achieve a zero-th order algorithm with $O(\delta^{-1}\epsilon^{-3})$ convergence and a first-order algorithm with $O(\delta^{-1}\epsilon^{-4})$ convergence to a ($\epsilon,\delta$)-Goldstein stationary point. On the other hand, yao2024constrained; lu2024first consider a different stationarity using $\epsilon$-KKT stationary, where lu2024first achieves a $O(\epsilon^{-7})$ convergence rate, and yao2024constrained achieves a $O(\epsilon^{-2})$ rate under a stronger assumption of access to projection operators.
Nonsmooth and nonconvex optimization. The nonsmooth and nonconvex structure of bilevel optimization with constraints makes its analysis closely related to nonsmooth nonconvex optimization. The best-known convergence result is given by zhang2020complexity, which establishes optimal convergence rates of $O(\delta^{-1}\epsilon^{-3})$ for the deterministic case and $O(\delta^{-1}\epsilon^{-4})$ for the stochastic case. Our result of $\tilde{O}(\delta^{-1}\epsilon^{-5})$ is one factor of $\epsilon$ away from the optimal stochastic rate, indicating potential room for improvement.
3 Problem Formulation and Penalty-Based Approximation
We consider the following linearly constrained bilevel optimization problem:
$\min_{x\in X}F(x)$ $=\mathbb{E}_{\xi}\left[f(x,y^{*}(x);\xi)\right]$ (2)
$\text{s.t.}\quad y^{*}(x)$ $\in\operatorname*{arg\,min}_{y\in\mathbb{R}^{m}:h(x,y)\leq\mathbf{0}}\mathbb{E}_{\zeta}\left[g(x,y;\zeta)\right]$
Here, $x\in\mathbb{R}^{n}$ denotes the upper-level (UL) decision variable constrained to a closed convex set $X\subseteq\mathbb{R}^{n}$, and $y\in\mathbb{R}^{m}$ is the lower-level (LL) variable. The UL and LL objective functions $f(x,y;\xi)$ and $g(x,y;\zeta)$ are stochastic, depending on random variables $\xi$ and $\zeta$, respectively, which model data or simulation noise. Expectations are taken with respect to the underlying distributions of these random variables.
The LL feasible region is defined by a set of $p$ linear inequality constraints:
$h(x,y):=\mathbf{A}x-\mathbf{B}y-\mathbf{b}\leq\mathbf{0},$ (3)
where $\mathbf{A}\in\mathbb{R}^{p\times n}$, $\mathbf{B}\in\mathbb{R}^{p\times m}$, and $\mathbf{b}\in\mathbb{R}^{p}$ are known matrices and vector. We assume the norm of the matrices are bounded by a given constant: $\|A\|\leq M_{\nabla h}$ and $\|B\|\leq M_{\nabla h}$, which also ensures that the Jacobian of the constraint function $h(x,y)$ is bounded.
Directly solving the stochastic bilevel problem is challenging due to the implicit dependence of $F(x)$ on $y^{*}(x)$ and the presence of noise in gradient and function evaluations.
3.1 Assumptions
We apply the following standard assumptions to our problem.
Assumption 3.1 (Smoothness and Strong Convexity).
We make the following assumptions on the objectives $f,g$, constraints $h$, and associated matrices:
  • (i)
    Upper-Level Objective $f$: The function $f(x,y)$ is $C_{f}$-smooth in $(x,y)$ (i.e., its gradient $\nabla f$ is $C_{f}$-Lipschitz continuous). The function $f(x,y)$ is also $L_{f}$-Lipschitz continuous in $(x,y)$.
    Note: $L_{f}$-Lipschitz continuity of $f$ implies that its gradient norm is bounded, i.e., $\|\nabla f(x,y)\|\leq L_{f}$.
  • (ii)
    Lower-Level Objective $g$: The function $g(x,y)$ is $C_{g}$-smooth in $(x,y)$ (i.e., its gradient $\nabla g$ is $C_{g}$-Lipschitz continuous). For each fixed $x\in X$, the function $g(x,\cdot)$ is $\mu_{g}$-strongly convex in $y$, with $\mu_{g}>0$. The gradient norm is bounded: $\|\nabla g(x,y)\|\leq L_{g}$. (Here, $L_{g}$ is a positive constant bounding the norm of $\nabla g$.)
  • (iii)
    Constraint Qualification (LICQ): The Linear Independence Constraint Qualification holds for the lower-level constraints at the optimal solution $y^{*}(x)$ for all $x\in X$. (Specifically, the Jacobian of the active constraints with respect to $y$, given by $\mathbf{B}$ restricted to its active rows, has full row rank.)
Under Assumption 3.1, the uniqueness of the LL solution $y^{*}(x)$ and the corresponding multipliers $\lambda^{*}(x)$ is guaranteed.
Assumption 3.2 (Global Lipschitz continuity of LL solution).
The optimal lower-level solution map $y^{*}(x)$ is globally $L_{y}$-Lipschitz continuous in $x$ on $X$. That is, there exists $L_{y}\geq 0$ such that $\|y^{*}(x)-y^{*}(x^{\prime})\|\leq L_{y}\|x-x^{\prime}\|$ for all $x,x^{\prime}\in X$.
This Lipschitz assumption is a standard assumption in bilevel optimization analysis to ensure convergence analysis (facchinei2003finite). Given the global Lipschitzness, the resulting UL objective $F(x)=f(x,y^{*}(x))$ is then $L_{F}$-Lipschitz continuous with $L_{F}\leq L_{f,x}+L_{f,y}L_{y}$.
We assume access only to noisy first-order information via stochastic oracles.
Assumption 3.3 (Stochastic Oracles).
Stochastic first-order oracles (SFOs) $\nabla\tilde{f}(x,y,\xi)$, $\nabla\tilde{g}(x,y;\zeta)$ are available, satisfying:
  • (i)
    Unbiasedness: $\mathbb{E}[\nabla\tilde{f}(x,y;\xi)]=\nabla f(x,y)$ and $\mathbb{E}[\nabla\tilde{g}(x,y;\zeta)]=\nabla g(x,y)$.
  • (ii)
    Bounded Variance: $\mathbb{E}[\|\nabla\tilde{f}-\nabla f\|^{2}]\leq\sigma^{2}$ and $\mathbb{E}[\|\nabla\tilde{g}-\nabla g\|^{2}]\leq\sigma^{2}$.
3.2 Goldstein Stationarity
Due to the potential nonsmoothness of $F(x)$, we target Goldstein stationarity, a robust concept for nonsmooth optimization.
Definition 3.1 (Goldstein Subdifferential (goldstein1977)).
For an $L_{F}$-Lipschitz function $F:\mathbb{R}^{n}\to\mathbb{R}$, $x\in\mathbb{R}^{n}$, $\delta>0$:
$\partial_{\delta}F(x):=\operatorname{conv}\left(\bigcup_{z\in B_{\delta}(x)}\partial F(z)\right)$
where $\partial F(z)$ is the Clarke subdifferential, and $B_{\delta}(x)$ is the $\delta$-ball around $x$.
Definition 3.2 ($(\delta,\epsilon)$-Goldstein Stationarity).
A point $x\in X$ is $(\delta,\epsilon)$-Goldstein stationary if
$dist(\mathbf{0},\partial_{\delta}F(x)+\mathcal{N}_{X}(x))\leq\epsilon,$
where $\mathcal{N}_{X}(x)$ is the normal cone to $X$ at $x$.
4 Error Analysis for Stochastic Hypergradient
We first control the effect of inexact dual variables on the penalty gradient and propagate this control to the shift of the penalized lower-level minimizer (Lemmas 4.1 and 4.2), which together yield an $O(\alpha)$ bias for the oracle (Lemma 4.3). We then bound the sampling variance by $O(1/N_{g})$ (Lemma 4.4); Theorem 4.1 consolidates these bounds, and Lemma 4.5 records the inner method cost $\tilde{O}(\alpha^{-2})$.
4.1 Stochastic Implementation
We compute the stochastic hypergradient oracle via a penalty formulation with smooth activation as
Algorithm 1 Stochastic Penalty-Based Hypergradient Oracle
1:Input: Point $x\in\mathbb{R}^{n}$, accuracy parameter $\alpha>0$, stochastic variance $\sigma^{2}$, batch size $N_{g}$
2:Set: $\alpha_{1}=\alpha^{-2}$, $\alpha_{2}=\alpha^{-4}$, $\delta=\alpha^{3}$
3: Compute $\tilde{y}^{*}(x)$ and $\tilde{\lambda}(x)$ by SGD such that $\|\tilde{y}^{*}(x)-y^{*}(x)\|\leq O(\delta)$ and $\|\tilde{\lambda}(x)-\lambda^{*}(x)\|\leq O(\delta)$
4: Define the smooth Lagrangian $L_{\tilde{\lambda},\alpha}(x,y)$ using the penalty Lagrangian and smooth activation function $\rho(x)$
5: Compute $\tilde{y}(x)=\operatorname*{arg\,min}_{y}L_{\tilde{\lambda},\alpha}(x,y)\|\leq\delta$ by SGD
6: Compute $\nabla\tilde{F}(x)=\frac{1}{N_{g}}\sum_{j=1}^{N_{g}}\nabla_{x}\tilde{L}_{\tilde{\lambda},\alpha}(x,\tilde{y}(x);\xi_{j})$ by $N_{g}$ independent samples
7:Output: $\nabla\tilde{F}(x)$
Remark 4.1 (Inner Loop Complexity).
With $g(x,\cdot)$ strongly convex and smooth, an SGD inner loop attains $O(\delta)$ accuracy in $O(\kappa_{g}\log(1/\delta))$ iterations. Setting $\delta=\Theta(\alpha^{3})$ and $\alpha=\Theta(\epsilon)$ yields $O(\kappa_{g}\log(1/\epsilon))$ iterations, so the inner workload grows only logarithmically in $1/\epsilon$.
Smooth Activation Function: To regularize constraint activation near the boundary, define $\rho_{i}(x)=\sigma_{h}(h_{i}(x,\tilde{y}^{*}(x)))\cdot\sigma_{\lambda}(\tilde{\lambda}_{i}(x))$ where:
$\sigma_{h}(z)=\begin{cases}0&\text{if }z<-\tau\delta\\ \frac{\tau\delta+z}{\tau\delta}&\text{if }-\tau\delta\leq z<0\\ 1&\text{if }z\geq 0\end{cases},\quad\sigma_{\lambda}(z)=\begin{cases}0&\text{if }z\leq 0\\ \frac{z}{\epsilon_{\lambda}}&\text{if }0<z<\epsilon_{\lambda}\\ 1&\text{if }z\geq\epsilon_{\lambda}\end{cases}$
with $\tau=\Theta(\delta)$ and $\epsilon_{\lambda}>0$ being small positive parameters.
The penalty function for hypergradient estimation is:
$L_{\tilde{\lambda},\alpha}(x,y)$ $=f(x,y)+\alpha_{1}\left(g(x,y)+(\tilde{\lambda}(x))^{T}h(x,y)\right.\left.-g(x,\tilde{y}^{*}(x))\right)+\frac{\alpha_{2}}{2}\sum_{i=1}^{p}\rho_{i}(x)\cdot h_{i}(x,y)^{2}$ (4)
where $\alpha_{1}=\alpha^{-2}$ and $\alpha_{2}=\alpha^{-4}$ for $\alpha>0$. The terms with $\tilde{\lambda}(x)$ and $\tilde{y}^{*}(x)$ promote KKT consistency and enforce constraints through a smoothed quadratic penalty.
The oracle outputs $\nabla\tilde{F}(x)$ with expectation $\mathbb{E}[\nabla\tilde{F}(x)]=\nabla_{x}L_{\tilde{\lambda},\alpha}(x,\tilde{y}(x))$. Its mean-squared error decomposes into bias and variance relative to $\nabla F(x)$:
$\mathbb{E}[\|\nabla\tilde{F}(x)-\nabla F(x)\|^{2}]$ $=\underbrace{\mathbb{E}[\|\nabla\tilde{F}(x)-\mathbb{E}[\nabla\tilde{F}(x)]\|^{2}]}_{\text{Variance}}+\underbrace{\|\mathbb{E}[\nabla\tilde{F}(x)]-\nabla F(x)\|^{2}}_{\text{Bias}^{2}}$ (5)
We first bound the effect of using $\tilde{\lambda}(x)$ in place of $\lambda^{*}(x)$ on the gradient of the penalty Lagrangian.
Lemma 4.1 (Lagrangian Gradient Approximation).
Assume $\|\tilde{\lambda}(x)-\lambda^{*}(x)\|\leq C_{\lambda}\delta$ and under Assumption 3.1 (iii), let $\alpha_{1}=\alpha^{-2}$, $\alpha_{2}=\alpha^{-4}$, and $\tau=\Theta(\delta)$. Then for fixed $(x,y)$:
$\|\nabla L_{\lambda^{*},\alpha}(x,y)-\nabla L_{\tilde{\lambda},\alpha}(x,y)\|\leq O(\alpha_{1}\delta+\alpha_{2}\delta).$
Proof sketch. Consider $\Delta\lambda:=\lambda^{*}(x)-\tilde{\lambda}(x)$ and decompose the gradient difference into a linear penalty part and a quadratic penalty part. For the linear term, $\|\nabla h\|$ is bounded by Assumption 3.1(iii) and $\|\Delta\lambda\|\leq C_{\lambda}\delta$, yielding $O(\alpha_{1}\delta)$. For the quadratic term, only near-active constraints contribute where $|h_{i}(x,\tilde{y}^{*}(x))|\leq\tau\delta$, and one obtains two pieces: $\alpha_{2}\sum_{i}(\rho_{i}^{*}-\tilde{\rho}_{i})h_{i}\nabla h_{i}$ and $\tfrac{\alpha_{2}}{2}\sum_{i}h_{i}^{2}\nabla(\rho_{i}^{*}-\tilde{\rho}_{i})$. Using $|h_{i}|=O(\delta)$ in these regions, $\|\nabla h_{i}\|$ bounded, and $\|\nabla(\rho_{i}^{*}-\tilde{\rho}_{i})\|=O(1/\delta)$, both pieces are $O(\alpha_{2}\delta)$. Combining yields $O(\alpha_{1}\delta+\alpha_{2}\delta)$.
Building on this result, we next bound the difference between exact solutions for true and approximate dual variables:
Lemma 4.2 (Solution Error).
Let $y^{*}_{\lambda,\alpha}(x):=\operatorname*{arg\,min}_{y}L_{\lambda,\alpha}(x,y)$ with $\alpha_{1}=\alpha^{-2}$ and $\alpha_{2}=\alpha^{-4}$.
Assume the target accuracy parameter $\alpha$ satisfies $\alpha\geq\frac{2C_{f}}{\mu_{g}}$ (where $C_{f}$ is the smoothness constant of $f$ and $\mu_{g}$ is the strong convexity constant of $g(x,\cdot)$ as per Assumption 3.1),
so that $L_{\lambda,\alpha}(x,y)$ is $\mu=\Omega(\alpha\mu_{g})$-strongly convex in $y$.
If the dual approximation satisfies $\|\tilde{\lambda}(x)-\lambda^{*}(x)\|\leq C_{\lambda}\delta$ and the gradient bound from Lemma 4.1 holds, then:
$\|y^{*}_{\lambda^{*},\alpha}(x)-y^{*}_{\tilde{\lambda},\alpha}(x)\|\leq\frac{C_{\text{sol}}}{\mu}(\alpha_{1}+\alpha_{2})\delta,$
where the constant $C_{\text{sol}}$ depends on $C_{\lambda}$ and $M_{\nabla h}$ (Assumption 3.1(iii) on $\|\nabla h\|$ bound).
Strong convexity together with Lemma 4.1 yields the stated solution bound.
With controlled approximation errors, we now derive a systematic bias bound.
4.2 Bias Analysis (Deterministic Error)
[…]
Lemma 4.3 (Hypergradient Bias Bound).
Let $\nabla_{x}L_{\lambda,\alpha}(x,y)$ denote the partial gradient of the penalty Lagrangian with respect to $x$. Assume it is $L_{H,y}$-Lipschitz in $y$ and $L_{H,\lambda}$-Lipschitz in $\lambda$. With $\alpha_{1}=\alpha^{-2}$, $\alpha_{2}=\alpha^{-4}$, choose $\delta=\Theta(\alpha^{3})$ and suppose $\|\tilde{y}(x)-y^{*}_{\tilde{\lambda},\alpha}(x)\|\leq\delta$ and $\|\tilde{\lambda}(x)-\lambda^{*}(x)\|\leq C_{\lambda}\delta$. If $L_{\lambda^{*},\alpha}(x,\cdot)$ is $\mu$-strongly convex with $\mu\geq c_{\mu}\alpha^{-2}$, then
$\|\mathbb{E}[\nabla\tilde{F}(x)]-\nabla F(x)\|\leq C_{\text{bias}}\alpha,$
where $C_{\text{bias}}$ depends only on $L_{H,y}$, $L_{H,\lambda}$, $C_{g}$, $C_{\lambda}$, $c_{\mu}$, and the penalty constant $C_{\text{pen}}$.
Proof sketch.
By the triangle inequality, write $\|\mathbb{E}[\nabla\tilde{F}(x)]-\nabla F(x)\|\leq T_{1}+T_{2}+T_{3}$, where $T_{1}=\|\nabla_{x}L_{\tilde{\lambda},\alpha}(x,\tilde{y}(x))-\nabla_{x}L_{\tilde{\lambda},\alpha}(x,y^{*}_{\tilde{\lambda},\alpha}(x))\|\leq L_{H,y}\,\delta=O(\alpha^{3})$ for $\delta=\Theta(\alpha^{3})$;
$T_{2}=\|\nabla_{x}L_{\tilde{\lambda},\alpha}(x,y^{*}_{\tilde{\lambda},\alpha}(x))-\nabla_{x}L_{\lambda^{*},\alpha}(x,y^{*}_{\lambda^{*},\alpha}(x))\|\leq L_{H,\lambda}C_{\lambda}\,\delta+L_{H,y}\,\|y^{*}_{\tilde{\lambda},\alpha}(x)-y^{*}_{\lambda^{*},\alpha}(x)\|$ and Lemma 4.2 gives $\|y^{*}_{\tilde{\lambda},\alpha}-y^{*}_{\lambda^{*},\alpha}\|\leq(C_{\text{sol}}/\mu)(\alpha_{1}+\alpha_{2})\delta$, so with $\alpha_{1}=\alpha^{-2}$, $\alpha_{2}=\alpha^{-4}$, $\delta=\Theta(\alpha^{3})$, $\mu=\Theta(\alpha^{-2})$ we obtain $T_{2}=O(\alpha)$;
$T_{3}=\|\nabla_{x}L_{\lambda^{*},\alpha}(x,y^{*}_{\lambda^{*},\alpha}(x))-\nabla F(x)\|=O(C_{\text{pen}}\alpha)$ by kornowski2024. Hence $\|\mathbb{E}[\nabla\tilde{F}(x)]-\nabla F(x)\|=O(\alpha)$. ∎
Thus the bias scales as $O(\alpha)$ when the inner accuracy is set to $\delta=\Theta(\alpha^{3})$.
4.3 Variance Analysis (Stochastic Error)
The variance, $\operatorname{Var}(\nabla\tilde{F}(x)\mid x,\tilde{\lambda},\tilde{y})=\mathbb{E}[\|\nabla\tilde{F}(x)-\mathbb{E}[\nabla\tilde{F}(x)]\|^{2}\mid x,\tilde{\lambda},\tilde{y}$], quantifies the error due to a finite batch size $N_{g}$ in estimating $\mathbb{E}[\nabla\tilde{F}(x)]$.
Lemma 4.4 (Variance Bound).
Under Assumption 3.3 (i)–(ii), let $\sigma^{2}$ be a uniform bound on
[…]
With a mini‑batch of $N_{g}$ i.i.d. samples in Algorithm 1, the conditional variance of the hypergradient estimate satisfies
$\operatorname{Var}\!\bigl(\nabla\tilde{F}(x)\,\big|\,x,\tilde{\lambda},\tilde{y}\bigr)\;\leq\;\frac{\sigma^{2}}{N_{g}}.$
Proof Sketch.
The hypergradient estimate $\nabla\tilde{F}(x)$ is the average of $N_{g}$ i.i.d. random vectors $G_{j}=\nabla_{x}\tilde{L}_{\tilde{\lambda},\alpha}(x,\tilde{y};\xi_{j})$.
By Assumption 3.3(ii), each term $G_{j}$ has conditional variance bounded by $\sigma^{2}$, i.e.,
$\operatorname{Var}(G_{j}\mid x,\tilde{\lambda},\tilde{y})\leq\sigma^{2}$. Since the samples $\xi_{j}$ are i.i.d., the terms $G_{j}$ are conditionally independent, and the conditional variance of their average is bounded accordingly. This follows standard mini-batch averaging analysis. ∎
4.4 Combined Error Bounds
We combine the bias and variance bounds to characterize the overall accuracy of the hypergradient oracle.
Theorem 4.1 (Accuracy of Stochastic Hypergradient).
Let $\nabla\tilde{F}(x)$ be the output of Algorithm 1 with penalty parameters $\alpha_{1}=\alpha^{-2},\alpha_{2}=\alpha^{-4}$, and inner accuracy $\delta=O(\alpha^{3})$. There exists a constant $C_{\text{bias}}$ such that:
$\mathbb{E}[\|\nabla\tilde{F}(x)-\nabla F(x)\|^{2}]\leq 2C_{\text{bias}}^{2}\alpha^{2}+\frac{2\sigma^{2}}{N_{g}}.$
Proof Sketch.
We apply the standard bias–variance decomposition. The bias term is bounded using Lemma 4.3, which shows that the expected output of the oracle approximates the true gradient up to an $O(\alpha)$ error. - The variance term is controlled using Lemma 4.4, which shows that averaging $N_{g}$ noisy gradients leads to variance bounded by $\sigma^{2}/N_{g}$.
Adding these two contributions and applying Jensen’s inequality yields the desired total error bound. ∎
Lemma 4.5 (Inner-loop Oracle Complexity).
Fix $\alpha>0$ and set $\alpha_{1}=\alpha^{-2}$, $\alpha_{2}=\alpha^{-4}$, $\delta=\Theta(\alpha^{3})$. Let $g(x,\cdot)$ be $\mu_{g}$-strongly convex and $C_{g}$-smooth, and the stochastic oracles of Assumption 3.3 have variance $\sigma^{2}$. Choose the mini-batch size $N_{g}=\sigma^{2}/\alpha^{2}$. Running Algorithm 1 with $\tilde{O}\!\bigl(\alpha^{-2}\bigr)$ stochastic first-order oracle (SFO) calls in its inner loops yields a stochastic inexact gradient $\nabla\tilde{F}(x)$ characterized by bias of $O(\alpha)$ and variance of $O(\alpha^{2})$.
5 Stochastic Bilevel Algorithm and Convergence Analysis
We now introduce the principal algorithm, F2CSA (Algorithm 2), which leverages the previously analyzed stochastic hypergradient oracle within a non-smooth, non-convex optimization framework.
We then present detailed convergence proofs that provide rigorous guarantees for identifying $(\delta,\epsilon)$-Goldstein stationary points.
Algorithm 2 Nonsmooth Nonconvex Algorithm with Inexact Stochastic Hypergradient Oracle
1:Input: Initialization $x_{0}\in\mathbb{R}^{n}$, clipping parameter $D>0$, step size $\eta>0$, Goldstein accuracy $\delta>0$, iteration budget $T\in\mathbb{N}$, inexact stochastic gradient oracle $\nabla\tilde{F}:\mathbb{R}^{n}\to\mathbb{R}^{n}$
2:Initialize: $\Delta_{1}=0$
3:for $t=1,\ldots,T$ do
4:  Sample $s_{t}\sim\text{Unif}[0,1]$
5:  $x_{t}=x_{t-1}+\Delta_{t}$, $z_{t}=x_{t-1}+s_{t}\Delta_{t}$
6:  Compute $g_{t}=\nabla\tilde{F}(z_{t})$ by running Algorithm 1 with $N_{g}=\Theta(\sigma^{2}/\alpha^{2})$ samples, so the inexact gradient has bias $O(\alpha)$ and variance $O(\alpha^{2})$.
7:  $\Delta_{t+1}=\text{clip}_{D}(\Delta_{t}-\eta g_{t})$ $\text{clip}_{D}(v):=\min\{1,\frac{D}{\|v\|}\}\cdot v$}
8:end for
[…]
10:for $k=1,\ldots,K$ do
11:  $x_{k}=\frac{1}{M}\sum_{m=1}^{M}z_{(k-1)M+m}$
12:end for
13:Output: $x_{\text{out}}\sim\text{Uniform}\{x_{1},\ldots,x_{K}\}$
Algorithm 2 provides an iterative framework leveraging our inexact stochastic hypergradient oracle. The method maintains a direction term $\Delta_{t}$, updated using a momentum-like step involving the oracle’s output $g_{t}=\nabla\tilde{F}(z_{t})$ and subsequently clipped to ensure $\|\Delta_{t}\|\leq D$. The output iterates $x_{k}$ are constructed by averaging sample points $z_{t}$ to approximate the Goldstein subdifferential goldstein1977; zhang2020complexity; davis2019stochastic.
Remark 5.1 (Integration with Stochastic Hypergradient Oracle).
Algorithm 2 utilizes Algorithm 1 as its gradient estimation subroutine. This integration needs careful parameter adjustments for accurate oracle estimates and manageable computation.
5.1 Convergence to Goldstein Stationarity
The following theorem establishes that with an inexact gradient oracle having bounded error, Algorithm 2 achieves Goldstein stationarity with a specified number of iterations.
Theorem 5.1 (Convergence with Stochastic Hypergradient Oracle).
[…]
  1. 1.
    Bias bound: $\|\mathbb{E}[\nabla\tilde{F}(x)]-\nabla F(x)\|\leq C_{\text{bias}}\alpha$
  2. 2.
    Variance bound: $\mathbb{E}[\|\nabla\tilde{F}(x)-\mathbb{E}[\nabla\tilde{F}(x)]\|^{2}]\leq\frac{\sigma^{2}}{N_{g}}$
[…]
Proof Sketch.
We first take $N_{g}=\Theta(\sigma^{2}/\alpha^{2})$ to get bias $O(\alpha)$ and variance $O(\alpha^{2})$.
Clip online gradient descent telescopes; error terms are $O(\eta)$ (stability) and $O(D\alpha)$ (stochastic).
Choose $M=\Theta(\epsilon^{-2})$ and $D=\Theta(\delta\epsilon^{2})$ so $\|z_{t}-x_{k}\|\leq MD\leq\delta$ and the block average lies in $\partial_{\delta}F(x_{k})$.
Set $\eta=\Theta(\delta\epsilon^{3})$ and run $T=O(((F(x_{0})-\inf F)L_{F}^{2})/(\delta\epsilon^{3}))$ to obtain $\mathbb{E}[\operatorname{dist}(\mathbf{0},\partial_{\delta}F(x_{\text{out}}))]\leq\epsilon+O(\alpha)$.
Finally, set $\alpha=\Theta(\epsilon)$ to conclude $\Theta(\epsilon)$ stationarity. ∎
Using Theorem 5.1, we can finally show the overall complexity of stochastic constrained bilevel optimization.
Theorem 5.2 (Complexity of solving stochastic constrained bilevel optimization).
The total stochastic first-order oracle (SFO) complexity is
$\begin{aligned} T\cdot N_{g}&=\Theta\left(\frac{F(x_{0})-\inf F}{\delta\epsilon^{3}}\right)\cdot\Theta\left(\frac{\sigma^{2}}{\epsilon^{2}}\right)=\Theta\left(\frac{(F(x_{0})-\inf F)\sigma^{2}}{\delta\epsilon^{5}}\right)\end{aligned}$ (6)
Including logarithmic factors from the inner loops, this becomes:
[…]
6 Experiments
To validate our theoretical analysis and assess the practical performance of the proposed F2CSA algorithm, we conduct experiments on synthetic bilevel optimization problems. We compare our method against SSIGD khanduri2023linearly and DSBLO khanduri2024doubly, both Hessian-based approaches by Khanduri et al. SSIGD uses an implicit gradient approach while DSBLO employs a doubly stochastic bilevel method. These comparisons highlight the computational advantages of our first-order approach over methods requiring second-order information.
6.1 Problem Setup
We evaluate our approach on toy bilevel problems with box constraints:
$\min\nolimits_{x\in\mathbb{R}^{d}}$ $f(x,y^{*}(x))\coloneqq\tfrac{1}{2}x^{\top}Q_{u}x+c_{u}^{\top}x+\tfrac{1}{2}y^{\top}Py+x^{\top}Py$ (8)
s.t. $y^{*}(x)\in\operatorname*{arg\,min}\nolimits_{y\in[-\mathbf{1},\mathbf{1}]}g(x,y)\coloneqq\tfrac{1}{2}y^{\top}Q_{l}y+c_{l}^{\top}y+x^{\top}y$ (9)
Parameters $Q_{u},Q_{l},P,c_{u},c_{l}$ are sampled from zero-mean Gaussians. Stochasticity is introduced by adding Gaussian noise $\mathcal{N}(0,\sigma^{2})$ to the quadratic terms during gradient evaluations with noise standard deviation $\sigma=0.01$. All algorithms use identical problem instances, initial points, random seeds, and the same lower-level solver to ensure fair evaluation. Step sizes are calibrated to be comparable across methods: SSIGD employs diminishing step sizes with $\beta=10^{-4}$, DSBLO uses adaptive step size selection, and F2CSA utilizes fixed step size $\eta=10^{-5}$, reflecting their different algorithmic structures.
6.2 Results and Analysis
Figure 1: Loss convergence trajectories for F2CSA, SSIGD, and DSBLO in dimension 50.
Figure 2: Computational cost scaling with problem dimension.
Convergence Performance
Figure 2 shows convergence trajectories on a 50-dimensional problem. All three methods converge to similar final loss values, with F2CSA maintaining stable convergence. The comparable performance of F2CSA to Hessian-based methods is consistent with our theoretical analysis in Lemma 4.3, which bounds the oracle error at $O(\alpha)$ bias and $O(1/N_{g})$ variance.
Computational Scalability
[…]
Key Insights
[…]
7 Conclusion and Future Work
We introduced a fully first-order framework for linearly constrained stochastic bilevel optimization and established the first finite-time guarantee to $(\delta,\epsilon)$-Goldstein stationarity using a smoothed penalty-based hypergradient oracle. Section 4 quantified the oracle’s error via an $O(\alpha)$ bias and $O(1/N_{g})$ variance, which, together with the inner-loop cost in Lemma 4.5, yielded the calibrated choice $N_{g}=\Theta(\sigma^{2}/\alpha^{2})$ and inner tolerance $\delta=\Theta(\alpha^{3})$. Section 5 integrated this oracle into a clipped nonsmooth algorithm attaining $\mathbb{E}[\operatorname{dist}(0,\partial_{\delta}F(x_{\text{out}}))]\leq\epsilon+O(\alpha)$ in $T=O(((F(x_{0})-\inf F)L_{F}^{2})/(\delta\epsilon^{3}))$ iterations; setting $\alpha=\Theta(\epsilon)$ implies the overall SFO complexity $\tilde{O}(\delta^{-1}\epsilon^{-5})$. Experiments corroborated the theory: F2CSA achieves comparable convergence performance to Hessian-based methods while scaling favorably in high dimensions, outperforming baselines in wall-clock time at large $d$ without sacrificing solution quality.
Two limitations are noteworthy. First, the rate is one factor of $\epsilon$ from the best-known stochastic nonsmooth complexity, suggesting headroom for variance reduction or momentum to close this gap. Second, our analysis hinges on LICQ, strong convexity of the lower level, and linear constraints; relaxing any of these raises technical challenges in bias control and stability of the penalized subproblem.
Promising directions include: (i) tighter oracle design with variance-reduced estimators or momentum (e.g., SPIDER-style or clipping-plus-momentum) to approach the optimal $\tilde{O}(\delta^{-1}\epsilon^{-4})$ dependence; (ii) structure-aware penalties that remain stable under weaker qualifications or partial degeneracy; (iii) specialized treatments of one-sided stochasticity (UL-only or LL-only), where improved batching can sharpen constants and exponents; and (iv) extending the smoothing-penalty machinery to certain nonlinear constraints while preserving first-order implementability and finite-time control. These steps would broaden practicality in meta-learning, RL, and large-scale ERM scenarios where bilevel structure and constraints are intrinsic.
Appendix A Appendix
See 4.1
Proof.
Define penalty Lagrangian:
$L_{\lambda,\alpha}(x,y)=f(x,y)+\alpha_{1}(g(x,y)+\lambda^{T}h(x,y)-g(x,\tilde{y}^{*}(x)))$ (10)
$\quad+\frac{\alpha_{2}}{2}\sum_{i=1}^{p}\rho_{i}(x)h_{i}(x,y)^{2}$ (11)
with activation $\rho_{i}(x)=\sigma_{h}(h_{i}(x,\tilde{y}^{*}(x)))\cdot\sigma_{\lambda}(\lambda_{i}(x))$ and dual error $\Delta\lambda=\lambda^{*}(x)-\tilde{\lambda}(x)$.
The gradient difference decomposes as:
$\nabla L_{\lambda^{*},\alpha}-\nabla L_{\tilde{\lambda}}=\underbrace{\alpha_{1}(\nabla h)^{T}\Delta\lambda}_{\text{Linear Term}}+\underbrace{\nabla\Delta_{Q}}_{\text{Quadratic Term}}$ (12)
where $\Delta_{Q}=\frac{\alpha_{2}}{2}\sum_{i=1}^{p}\Delta\rho_{i}(x)h_{i}(x,y)^{2}$ with $\Delta\rho_{i}(x)=\rho_{i}^{*}(x)-\tilde{\rho}_{i}(x)$.
Linear Term in (12):
From $h(x,y)=\mathbf{A}x-\mathbf{B}y-\mathbf{b}$, we have:
$\|\nabla h\|\leq\|\mathbf{A}\|+\|\mathbf{B}\|\leq 2M_{AB}$ (13)
Using (13) and $\|\Delta\lambda\|\leq C_{\lambda}\delta$:
$\|\alpha_{1}(\nabla h)^{T}\Delta\lambda\|\leq\alpha_{1}\cdot 2M_{AB}\cdot C_{\lambda}\delta=O(\alpha_{1}\delta)$ (14)
Quadratic Term in (12):
The quadratic gradient expands to:
$\nabla\Delta_{Q}=\underbrace{\alpha_{2}\sum_{i=1}^{p}\Delta\rho_{i}h_{i}\nabla h_{i}}_{\text{$(T_{1})$}}+\underbrace{\frac{\alpha_{2}}{2}\sum_{i=1}^{p}h_{i}^{2}\nabla\Delta\rho_{i}}_{\text{T2}}$ (15)
(T1): $\Delta\rho_{i}\neq 0$ only for near-active constraints where $|h_{i}(x,\tilde{y}^{*}(x))|\leq\tau\delta=O(\delta)$.
For constraint values:
$h_{i}(x,y)-h_{i}(x,\tilde{y}^{*}(x))=\mathbf{B}_{i}(\tilde{y}^{*}(x)-y)$ (16)
$\implies|h_{i}(x,y)-h_{i}(x,\tilde{y}^{*}(x))|\leq M_{AB}\delta$ (17)
Using (17): $|h_{i}(x,y)|\leq O(\delta)+M_{AB}\delta=O(\delta)$ for near-active constraints.
With $|\Delta\rho_{i}|\leq 1$ and $\|\nabla h_{i}\|\leq 2M_{AB}$:
$\|\Delta\rho_{i}h_{i}\nabla h_{i}\|\leq O(\delta)\implies\|\text{T1}\|\leq\alpha_{2}p\cdot O(\delta)=O(\alpha_{2}\delta)$ (18)
$(T_{2})$: $\|\nabla\Delta\rho_{i}\|=O(1/\delta)$ only when $|h_{i}(x,\tilde{y}^{*}(x))|=O(\delta)$.
In these regions: $|h_{i}(x,y)|=O(\delta)$ so:
$h_{i}^{2}\cdot\|\nabla\Delta\rho_{i}\|=O(\delta)^{2}\cdot O(1/\delta)=O(\delta)$ (19)
Using (19):
$\|\text{T2}\|\leq\frac{\alpha_{2}}{2}p\cdot O(\delta)=O(\alpha_{2}\delta)$ (20)
From (14), (18), and (20):
$\|\nabla L_{\lambda^{*},\alpha}-\nabla L_{\tilde{\lambda}}\|\leq O(\alpha_{1}\delta)+O(\alpha_{2}\delta)+O(\alpha_{2}\delta)=O(\alpha_{1}\delta+\alpha_{2}\delta)$ (21)
See 4.2
Proof.
For brevity, let $y^{*}_{\lambda^{*},\alpha}(x)=y^{*}_{\lambda^{*}}$ and $y^{*}_{\tilde{\lambda},\alpha}(x)=y^{*}_{\tilde{\lambda}}$. From the definition of these minimizers, we have the first-order optimality conditions:
$\nabla_{y}L_{\lambda^{*},\alpha}(x,y^{*}_{\lambda^{*}})$ $=0$ (22)
[…]
The lemma assumes that $L_{\lambda,\alpha}(x,y)$ is $\mu$-strongly convex in $y$ (for relevant $\lambda$, including $\lambda^{*}$). Thus, $L_{\lambda^{*},\alpha}(x,\cdot)$ is $\mu$-strongly convex. A standard property of a $\mu$-strongly convex function $\phi(y)$ with minimizer $y_{1}^{*}$ is that for any $y_{2}$:
$\mu\|y_{1}^{*}-y_{2}\|^{2}\leq\langle\nabla_{y}\phi(y_{1}^{*})-\nabla_{y}\phi(y_{2}),y_{1}^{*}-y_{2}\rangle$
Applying this with $\phi(y)=L_{\lambda^{*},\alpha}(x,y)$, $y_{1}^{*}=y^{*}_{\lambda^{*}}$, and $y_{2}=y^{*}_{\tilde{\lambda}}$:
$\mu\|y^{*}_{\lambda^{*}}-y^{*}_{\tilde{\lambda}}\|^{2}\leq\langle\nabla_{y}L_{\lambda^{*},\alpha}(x,y^{*}_{\lambda^{*}})-\nabla_{y}L_{\lambda^{*},\alpha}(x,y^{*}_{\tilde{\lambda}}),y^{*}_{\lambda^{*}}-y^{*}_{\tilde{\lambda}}\rangle$
Using the optimality condition from Eq. (22), $\nabla_{y}L_{\lambda^{*},\alpha}(x,y^{*}_{\lambda^{*}})=0$, this simplifies to:
$\mu\|y^{*}_{\lambda^{*}}-y^{*}_{\tilde{\lambda}}\|^{2}\leq\langle-\nabla_{y}L_{\lambda^{*},\alpha}(x,y^{*}_{\tilde{\lambda}}),y^{*}_{\lambda^{*}}-y^{*}_{\tilde{\lambda}}\rangle$
[…]
$\mu\|y^{*}_{\lambda^{*}}-y^{*}_{\tilde{\lambda}}\|^{2}$ $\leq\langle\nabla_{y}L_{\tilde{\lambda},\alpha}(x,y^{*}_{\tilde{\lambda}})-\nabla_{y}L_{\lambda^{*},\alpha}(x,y^{*}_{\tilde{\lambda}}),y^{*}_{\lambda^{*}}-y^{*}_{\tilde{\lambda}}\rangle$
$\quad\textit{(since }\nabla_{y}L_{\tilde{\lambda},\alpha}(x,y^{*}_{\tilde{\lambda}})=0\textit{)}$
Applying the Cauchy-Schwarz inequality:
$\mu\|y^{*}_{\lambda^{*}}-y^{*}_{\tilde{\lambda}}\|^{2}\leq\|\nabla_{y}L_{\tilde{\lambda},\alpha}(x,y^{*}_{\tilde{\lambda}})-\nabla_{y}L_{\lambda^{*},\alpha}(x,y^{*}_{\tilde{\lambda}})\|\cdot\|y^{*}_{\lambda^{*}}-y^{*}_{\tilde{\lambda}}\|$
If $y^{*}_{\lambda^{*}}\neq y^{*}_{\tilde{\lambda}}$, we can divide by $\|y^{*}_{\lambda^{*}}-y^{*}_{\tilde{\lambda}}\|$:
[…]
Lemma 4.1 states that for any fixed $(x,y)$, $\|\nabla L_{\lambda^{*},\alpha}(x,y)-\nabla L_{\tilde{\lambda},\alpha}(x,y)\|\leq O(\alpha_{1}\delta+\alpha_{2}\delta)$. This implies there exists a constant, which we identify with $C_{\text{sol}}$ from the lemma statement (where $C_{\text{sol}}$ depends on $C_{\lambda}$ and $M_{\nabla h}$), such that:
$\|\nabla L_{\lambda^{*},\alpha}(x,y^{*}_{\tilde{\lambda}})-\nabla L_{\tilde{\lambda},\alpha}(x,y^{*}_{\tilde{\lambda}})\|\leq C_{\text{sol}}(\alpha_{1}+\alpha_{2})\delta$
Substituting this into the inequality above:
$\mu\|y^{*}_{\lambda^{*}}-y^{*}_{\tilde{\lambda}}\|\leq C_{\text{sol}}(\alpha_{1}+\alpha_{2})\delta$
Dividing by $\mu$ (which is positive as $\mu=\Omega(\alpha\mu_{g})$ and $\mu_{g}>0,\alpha>0$) yields the result:
$\|y^{*}_{\lambda^{*},\alpha}(x)-y^{*}_{\tilde{\lambda},\alpha}(x)\|\leq\frac{C_{\text{sol}}}{\mu}(\alpha_{1}+\alpha_{2})\delta$
If $y^{*}_{\lambda^{*}}=y^{*}_{\tilde{\lambda}}$, the inequality holds trivially. This completes the proof. ∎
See 4.3
Proof.
The quantity to bound is the bias $\|\mathbb{E}[\nabla\tilde{F}(x)]-\nabla F(x)\|=\|\nabla_{x}L_{\tilde{\lambda},\alpha}(x,\tilde{y}(x))-\nabla F(x)\|$. We decompose this error into three parts using the triangle inequality:
$\|\nabla_{x}L_{\tilde{\lambda},\alpha}(x,\tilde{y}(x))-\nabla F(x)\|$ $\leq\underbrace{\|\nabla_{x}L_{\tilde{\lambda},\alpha}(x,\tilde{y}(x))-\nabla_{x}L_{\tilde{\lambda},\alpha}(x,y^{*}_{\tilde{\lambda},\alpha}(x))\|}_{T_{1}}$ (24)
$\quad+\underbrace{\|\nabla_{x}L_{\tilde{\lambda},\alpha}(x,y^{*}_{\tilde{\lambda},\alpha}(x))-\nabla_{x}L_{\lambda^{*},\alpha}(x,y^{*}_{\lambda^{*},\alpha}(x))\|}_{T_{2}}$ (25)
$\quad+\underbrace{\|\nabla_{x}L_{\lambda^{*},\alpha}(x,y^{*}_{\lambda^{*},\alpha}(x))-\nabla F(x)\|}_{T_{3}}$ (26)
$(T_{1})$: This term bounds the error from the inexact minimization of $L_{\tilde{\lambda},\alpha}(x,\cdot)$. Using the $L_{H,y}$-Lipschitz continuity of $\nabla_{x}L_{\tilde{\lambda},\alpha}(x,y)$ with respect to $y$ (as assumed in the lemma statement) and the condition $\|\tilde{y}(x)-y^{*}_{\tilde{\lambda},\alpha}(x)\|\leq\delta$ (from the lemma statement, where $\delta=\Theta(\alpha^{3})$):
$T_{1}\leq L_{H,y}\|\tilde{y}(x)-y^{*}_{\tilde{\lambda},\alpha}(x)\|\leq L_{H,y}\delta=O(\delta).$ (27)
$(T_{2})$: This term bounds the error from using the approximate dual $\tilde{\lambda}(x)$ instead of the true dual $\lambda^{*}(x)$ in defining the penalty Lagrangian and its minimizer. Using the triangle inequality:
$T_{2}$ $\leq\|\nabla_{x}L_{\tilde{\lambda},\alpha}(x,y^{*}_{\tilde{\lambda},\alpha}(x))-\nabla_{x}L_{\tilde{\lambda},\alpha}(x,y^{*}_{\lambda^{*},\alpha}(x))\|$
$\quad+\|\nabla_{x}L_{\tilde{\lambda},\alpha}(x,y^{*}_{\lambda^{*},\alpha}(x))-\nabla_{x}L_{\lambda^{*},\alpha}(x,y^{*}_{\lambda^{*},\alpha}(x))\|.$ (28)
The first part of the sum is bounded by $L_{H,y}\|y^{*}_{\tilde{\lambda},\alpha}(x)-y^{*}_{\lambda^{*},\alpha}(x)\|$ (using the assumed $L_{H,y}$-Lipschitz continuity of $\nabla_{x}L_{\tilde{\lambda},\alpha}(x,y)$ w.r.t $y$). The second part is bounded by $L_{H,\lambda}\|\tilde{\lambda}(x)-\lambda^{*}(x)\|$ (using the assumed $L_{H,\lambda}$-Lipschitz continuity of $\nabla_{x}L_{\cdot,\alpha}(x,y^{*}_{\lambda^{*},\alpha}(x))$ w.r.t the dual variable). Invoking Lemma 4.2 for $\|y^{*}_{\tilde{\lambda},\alpha}(x)-y^{*}_{\lambda^{*},\alpha}(x)\|\leq\frac{C_{\text{sol}}}{\mu}(\alpha_{1}+\alpha_{2})\delta$, and using the condition $\|\tilde{\lambda}(x)-\lambda^{*}(x)\|\leq C_{\lambda}\delta$ (from Assumption 3.3, with $\delta=\Theta(\alpha^{3})$ as per this lemma’s setup):
$T_{2}$ $\leq L_{H,y}\cdot\frac{C_{\text{sol}}}{\mu}(\alpha_{1}+\alpha_{2})\delta+L_{H,\lambda}\cdot C_{\lambda}\delta$ (29)
$=O\left(\frac{(\alpha_{1}+\alpha_{2})\delta}{\mu}\right)+O(\delta).$ (30)
Given $\alpha_{1}=\alpha^{-2}$, $\alpha_{2}=\alpha^{-4}$, so $(\alpha_{1}+\alpha_{2})=O(\alpha^{-4})$. With $\delta=\Theta(\alpha^{3})$ and $\mu=\Theta(\alpha^{-2})$ (from the lemma condition $\mu\geq c_{\mu}\alpha^{-2}$), the first term is $O\left(\frac{\alpha^{-4}\alpha^{3}}{\alpha^{-2}}\right)=O(\alpha)$. The second term $O(\delta)$ is $O(\alpha^{3})$. Thus, $T_{2}=O(\alpha)$.
$(T_{3})$: This term measures the inherent approximation error of the idealized penalty method (using true $\lambda^{*}$ and exact minimization $y^{*}_{\lambda^{*},\alpha}(x)$) with respect to the true hypergradient $\nabla F(x)$. As per the lemma’s setup, this is bounded by:
$T_{3}=\|\nabla_{x}L_{\lambda^{*},\alpha}(x,y^{*}_{\lambda^{*},\alpha}(x))-\nabla F(x)\|\leq C_{\text{pen}}\alpha,$ (31)
for some problem-dependent constant $C_{\text{pen}}$.
Combining Terms: Summing the bounds for $T_{1},T_{2},$ and $T_{3}$, with $\delta=\Theta(\alpha^{3})$:
$\|\nabla_{x}L_{\tilde{\lambda},\alpha}(x,\tilde{y}(x))-\nabla F(x)\|$ $\leq O(\delta)+O(\alpha)+O(C_{\text{pen}}\alpha)$ (32)
$=O(\alpha^{3})+O(\alpha)+O(\alpha)=O(\alpha).$ (33)
Since $\mathbb{E}[\nabla\tilde{F}(x)]=\nabla_{x}L_{\tilde{\lambda},\alpha}(x,\tilde{y}(x))$, we conclude that $\|\mathbb{E}[\nabla\tilde{F}(x)]-\nabla F(x)\|\leq C_{\text{bias}}\alpha$. The conditions $\delta=\Theta(\alpha^{3})$ and $\mu=\Theta(\alpha^{-2})$ ensure that all error components are either $O(\alpha)$ or of a smaller order. ∎
See 4.4
Proof.
Let $G_{j}=\nabla_{x}\tilde{L}_{\tilde{\lambda},\alpha}(x,\tilde{y};\xi_{j})$ for $j=1,\dots,N_{g}$. Conditional on $x,\tilde{\lambda},\tilde{y}$, these $G_{j}$ are i.i.d. random vectors with mean $\mathbb{E}[G_{j}\mid x,\tilde{\lambda},\tilde{y}]=\nabla_{x}L_{\tilde{\lambda},\alpha}(x,\tilde{y}(x))=\mathbb{E}[\nabla\tilde{F}(x)\mid x,\tilde{\lambda},\tilde{y}$].
The conditional variance of the averaged estimator is:
[…]
Since the $G_{j}$ are independent conditional on $x,\tilde{\lambda},\tilde{y}$, we have:
[…]
By our Assumption 3.3(ii), $\operatorname{Var}(G_{j}\mid x,\tilde{\lambda},\tilde{y})\leq\sigma^{2}$ for all $j$. Therefore:
[…]
Thus, the variance of the hypergradient estimator is bounded by $\frac{\sigma^{2}}{N_{g}}$. ∎
See 4.1
Proof.
i) Using the bias-variance decomposition and properties of conditional expectation:
$\mathbb{E}[\|\nabla\tilde{F}(x)-\nabla F(x)\|^{2}]$ $=\mathbb{E}[\|\nabla\tilde{F}(x)-\mathbb{E}[\nabla\tilde{F}(x)]+\mathbb{E}[\nabla\tilde{F}(x)]-\nabla F(x)\|^{2}]$ (37)
By the inequality $\|a+b\|^{2}\leq 2\|a\|^{2}+2\|b\|^{2}$:
$\mathbb{E}[\|\nabla\tilde{F}(x)-\nabla F(x)\|^{2}]$ $\leq 2\mathbb{E}[\|\nabla\tilde{F}(x)-\mathbb{E}[\nabla\tilde{F}(x)]\|^{2}]+2\|\mathbb{E}[\nabla\tilde{F}(x)]-\nabla F(x)\|^{2}$ (38)
The first term is the expected conditional variance:
$\mathbb{E}[\|\nabla\tilde{F}(x)-\mathbb{E}[\nabla\tilde{F}(x)]\|^{2}]$ $=\mathbb{E}[\mathbb{E}[\|\nabla\tilde{F}(x)-\mathbb{E}[\nabla\tilde{F}(x)]\|^{2}\mid x,\tilde{\lambda},\tilde{y}$]] (39)
[…]
From Lemma 4.4, we know that $\operatorname{Var}(\nabla\tilde{F}(x)\mid x,\tilde{\lambda},\tilde{y})\leq\frac{\sigma^{2}}{N_{g}}$. Therefore:
$\mathbb{E}[\|\nabla\tilde{F}(x)-\mathbb{E}[\nabla\tilde{F}(x)]\|^{2}]\leq\frac{\sigma^{2}}{N_{g}}$ (41)
The second term is the squared bias, which from Lemma 4.3 is bounded by:
$\|\mathbb{E}[\nabla\tilde{F}(x)]-\nabla F(x)\|^{2}\leq(C_{\text{bias}}\alpha)^{2}=C_{\text{bias}}^{2}\alpha^{2}$ (42)
Combining these bounds:
$\mathbb{E}[\|\nabla\tilde{F}(x)-\nabla F(x)\|^{2}]$ $\leq 2\cdot\frac{\sigma^{2}}{N_{g}}+2\cdot C_{\text{bias}}^{2}\alpha^{2}$ (43)
$=2C_{\text{bias}}^{2}\alpha^{2}+\frac{2\sigma^{2}}{N_{g}}$ (44)
See 4.5
Proof.
We count the stochastic-gradient oracle calls made in one execution of Algorithm 1. The inner tolerance is $\delta=\Theta(\alpha^{3})$.
[…]
$t_{1}=O\!\bigl(\kappa_{g}\log(1/\delta)\bigr)=O\!\Bigl(\tfrac{C_{g}}{\mu_{g}}\log\tfrac{1}{\delta}\Bigr)$
oracle calls give $\|\,\tilde{y}^{*}-y^{*}\|,\;\|\,\tilde{\lambda}^{*}-\lambda^{*}\|\leq\delta$.
C2. Penalty minimisation (y~)($\tilde${y}): With $\alpha_{1}=\alpha^{-2}$ and $\alpha_{2}=\alpha^{-4}$ we analyze $L_{\tilde{\lambda}^{*},\alpha}(x,\cdot)$:
  • Strong convexity. The term $\alpha_{1}g$ contributes $\alpha_{1}\mu_{g}$; the smooth term $f$ can subtract at most $C_{f}$ curvature. For sufficiently small $\alpha$, $\mu_{\mathrm{pen}}\geq\alpha_{1}\mu_{g}/2$.
  • Smoothness. Because each $h_{i}$ is affine in $y$, the quadratic penalty has Hessian bounded by $\alpha_{2}\|B\|^{2}$, so $L_{\mathrm{pen}}=\Theta(\alpha_{2})$
Therefore the condition number is
$\kappa_{\mathrm{pen}}=\frac{L_{\mathrm{pen}}}{\mu_{\mathrm{pen}}}=\Theta\!\bigl(\alpha^{-2}/\mu_{g}\bigr).$
[…]
C3. Total inner cost: Summing $t_{1}$ and $t_{2}$ and adding the mini-batch evaluations:
$\text{cost}(x)=O\!\Bigl(\bigl(\tfrac{C_{g}}{\mu_{g}}+\tfrac{\alpha^{-2}}{\mu_{g}}\bigr)\log\tfrac{1}{\delta}\Bigr)\;+\;N_{g}.$
[…]
$\text{cost}(x)=\tilde{O}\!\bigl(\alpha^{-2}/\mu_{g}\bigr)+N_{g}.$
Using Lemma 4.4, $N_{g}$ should satisfy $\sigma/\sqrt{N_{g}}\;\asymp\;\alpha$, hence $N_{g}=\Theta(\sigma^{2}/\alpha^{2})$. Plugging in, $\text{cost}(x)=\tilde{O}(\alpha^{-2})$ ( constants depending on $\mu_{g}$ and $\sigma^{2}$ are absorbed).
With this batch size, $\mathbb{E}\bigl\|\tilde{\nabla}F(x)-\nabla F(x)\bigr\|\leq O(\alpha)+\sigma/\sqrt{N_{g}}=O(\alpha),$ so the oracle outputs an $\alpha$-accurate hyper-gradient.
Set $\alpha=\Theta(\varepsilon)$ for outer-loop tolerance $\varepsilon$; the inner cost becomes $\tilde{O}(\varepsilon^{-2})$
See 5.1
Proof.
For any $t\in[T]$, since $x_{t}=x_{t-1}+\Delta_{t}$, we have by the fundamental theorem of calculus:
$F(x_{t})-F(x_{t-1})$ $=\int_{0}^{1}\langle\nabla F(x_{t-1}+s\Delta_{t}),\Delta_{t}\rangle ds$ (45)
$=\mathbb{E}_{s_{t}\sim\text{Unif}[0,1]}[\langle\nabla F(x_{t-1}+s_{t}\Delta_{t}),\Delta_{t}\rangle]$ (46)
$=\mathbb{E}[\langle\nabla F(z_{t}),\Delta_{t}\rangle]$ (47)
where equation (47) follows from our algorithm’s definition of $z_{t}=x_{t-1}+s_{t}\Delta_{t}$. Summing over $t\in[T]=[K\times M]$:
$\inf F\leq F(x_{T})$ $=F(x_{0})+\sum_{t=1}^{T}\mathbb{E}[\langle\nabla F(z_{t}),\Delta_{t}\rangle]$ (48)
$=F(x_{0})+\underbrace{\sum_{k=1}^{K}\sum_{m=1}^{M}\mathbb{E}[\langle\nabla F(z_{(k-1)M+m}),\Delta_{(k-1)M+m}-u_{k}\rangle]}_{\text{regret of online gradient descent}}$
$\quad+\underbrace{\sum_{k=1}^{K}\sum_{m=1}^{M}\mathbb{E}[\langle\nabla F(z_{(k-1)M+m}),u_{k}\rangle]}_{\text{Gradient norm}}$ (49)
where we’ve added and subtracted $\langle\nabla F(z_{t}),u_{k}\rangle$ in (49) for any sequence of reference points $u_{1},\ldots,u_{K}\in\mathbb{R}^{d}$ satisfying $\|u_{i}\|\leq D$ for all $i$.
The first double sum represents the regret of online gradient descent with stochastic gradients. For any $t\in[T]$:
$\|\Delta_{t+1}-u_{k}\|^{2}$ $=\|\text{clip}_{D}(\Delta_{t}-\eta\tilde{g}_{t})-u_{k}\|^{2}$ (50)
$\leq\|\Delta_{t}-\eta\tilde{g}_{t}-u_{k}\|^{2}$ (51)
$=\|\Delta_{t}-u_{k}\|^{2}+\eta^{2}\|\tilde{g}_{t}\|^{2}-2\eta\langle\Delta_{t}-u_{k},\tilde{g}_{t}\rangle$ (52)
where (51) follows since projection onto a convex set decreases distance. Rearranging (52):
$\langle\tilde{g}_{t},\Delta_{t}-u_{k}\rangle\leq\frac{\|\Delta_{t}-u_{k}\|^{2}-\|\Delta_{t+1}-u_{k}\|^{2}}{2\eta}+\frac{\eta\|\tilde{g}_{t}\|^{2}}{2}$ (53)
Now, we decompose the key inner product using the bias-variance structure of our stochastic gradient oracle:
$\mathbb{E}[\langle\nabla F(z_{t}),\Delta_{t}-u_{k}\rangle]$ $=\mathbb{E}[\langle\tilde{g}_{t},\Delta_{t}-u_{k}\rangle]+\mathbb{E}[\langle\nabla F(z_{t})-\tilde{g}_{t},\Delta_{t}-u_{k}\rangle]$ (54)
First term in eq. 54:
For the first term in (54), using inequality (53):
$\mathbb{E}[\langle\tilde{g}_{t},\Delta_{t}-u_{k}\rangle]$ $\leq\mathbb{E}\left[\frac{\|\Delta_{t}-u_{k}\|^{2}-\|\Delta_{t+1}-u_{k}\|^{2}}{2\eta}+\frac{\eta\|\tilde{g}_{t}\|^{2}}{2}\right]$ (55)
For the expected squared norm in (55), using the bias-variance decomposition and the $L$-Lipschitz property of $F$:
$\mathbb{E}[\|\tilde{g}_{t}\|^{2}]$ $=\mathbb{E}[\|\mathbb{E}[\tilde{g}_{t}|z_{t}]+(\tilde{g}_{t}-\mathbb{E}[\tilde{g}_{t}|z_{t}])\|^{2}]$ (56)
$\leq\mathbb{E}[\|\mathbb{E}[\tilde{g}_{t}|z_{t}]\|^{2}]+\mathbb{E}[\|\tilde{g}_{t}-\mathbb{E}[\tilde{g}_{t}|z_{t}]\|^{2}]$ (57)
$\leq L^{2}+\frac{\sigma^{2}}{N_{g}}$ (58)
[…]
where (57) follows from the orthogonality of bias and variance terms. Therefore, we have:
$\mathbb{E}[\langle\tilde{g}_{t},\Delta_{t}-u_{k}\rangle]$ $\leq\mathbb{E}\left[\frac{\|\Delta_{t}-u_{k}\|^{2}-\|\Delta_{t+1}-u_{k}\|^{2}}{2\eta}\right]+O(\eta)$ (60)
Second term in eq. 54:
based on Cauchy-Schwarz inequality and noting that both $\|\Delta_{t}\|\leq D$ and $\|u_{k}\|\leq D$ by construction, we have:
$\mathbb{E}[\langle\nabla F(z_{t})-\tilde{g}_{t},\Delta_{t}-u_{k}\rangle]$ $\leq\mathbb{E}[\|\nabla F(z_{t})-\tilde{g}_{t}\|\cdot\|\Delta_{t}-u_{k}\|]$ (61)
$\leq 2D\cdot\mathbb{E}[\|\nabla F(z_{t})-\tilde{g}_{t}\|]$ (62)
By triangle inequality and the properties of our stochastic oracle:
$\mathbb{E}[\|\nabla F(z_{t})-\tilde{g}_{t}\|]$ $\leq\mathbb{E}[\|\nabla F(z_{t})-\mathbb{E}[\tilde{g}_{t}|z_{t}]\|]+\mathbb{E}[\|\mathbb{E}[\tilde{g}_{t}|z_{t}]-\tilde{g}_{t}\|]$ (63)
$\leq C_{\text{bias}}\alpha+\mathbb{E}[\|\mathbb{E}[\tilde{g}_{t}|z_{t}]-\tilde{g}_{t}\|]$ (64)
For the variance term in (64), by Jensen’s inequality:
$\mathbb{E}[\|\mathbb{E}[\tilde{g}_{t}|z_{t}]-\tilde{g}_{t}\|]$ $\leq\sqrt{\mathbb{E}[\|\mathbb{E}[\tilde{g}_{t}|z_{t}]-\tilde{g}_{t}\|^{2}]}$ (65)
$=\sqrt{\mathbb{E}[\mathbb{E}[\|\mathbb{E}[\tilde{g}_{t}|z_{t}]-\tilde{g}_{t}\|^{2}|z_{t}]]}$ (66)
$=\sqrt{\mathbb{E}[\operatorname{Var}(\tilde{g}_{t}|z_{t})]}\leq\frac{\sigma}{\sqrt{N_{g}}}$ (67)
where (66) follows from the tower property of conditional expectation, and (67) uses our variance bound assumption. When $N_{g}=\Theta(\frac{\sigma^{2}}{\alpha^{2}})$ ensures:
$\mathbb{E}[\|\nabla F(z_{t})-\tilde{g}_{t}\|]\leq C_{\text{bias}}\alpha+\frac{\sigma}{\sqrt{N_{g}}}=C_{\text{bias}}\alpha+O(\alpha)=O(\alpha)$ (68)
Therefore, combining (62) and (68), we can bound the second term in eq. 54 by:
$\mathbb{E}[\langle\nabla F(z_{t})-\tilde{g}_{t},\Delta_{t}-u_{k}\rangle]\leq 2D\cdot O(\alpha)=O(D\alpha)$ (69)
Putting together first term and second term
: Combining (60) and (69):
$\mathbb{E}[\langle\nabla F(z_{t}),\Delta_{t}-u_{k}\rangle]$ $\leq\mathbb{E}\left[\frac{\|\Delta_{t}-u_{k}\|^{2}-\|\Delta_{t+1}-u_{k}\|^{2}}{2\eta}\right]+O(\eta)+O(D\alpha)$ (70)
Summing (70) over t=(k1)M+mt=(k-1)M+m with $m=1,\ldots,M$ for a fixed $k$:
$\sum_{m=1}^{M}\mathbb{E}[\langle\nabla F(z_{(k-1)M+m}),\Delta_{(k-1)M+m}-u_{k}\rangle]$
$\leq$ $\sum_{m=1}^{M}\mathbb{E}\left[\frac{\|\Delta_{(k-1)M+m}-u_{k}\|^{2}-\|\Delta_{(k-1)M+m+1}-u_{k}\|^{2}}{2\eta}\right]+\sum_{m=1}^{M}O(\eta)+\sum_{m=1}^{M}O(D\alpha)$ (71)
$\leq$ $\frac{\mathbb{E}[\|\Delta_{(k-1)M+1}-u_{k}\|^{2}-\|\Delta_{(k-1)M+M+1}-u_{k}\|^{2}]}{2\eta}+O(M\eta)+O(MD\alpha)$ (72)
Since $\|\Delta_{t}\|\leq D$ and $\|u_{k}\|\leq D$, we have $\|\Delta_{t}-u_{k}\|\leq 2D~\forall t$. Therefore, we can further bound (72) by:
$\frac{\mathbb{E}[\|\Delta_{(k-1)M+1}-u_{k}\|^{2}-\|\Delta_{(k-1)M+M+1}-u_{k}\|^{2}]}{2\eta}+O(M\eta)+O(MD\alpha)$ (73)
$\leq$ $\frac{4D^{2}}{2\eta}+O(M\eta)+O(MD\alpha)$
$=$ $O(\frac{D^{2}}{\eta}+M\eta+MD\alpha)$ (74)
Since this inequality holds for all $\eta\in\mathbb{R}_{+}$, we can choose $\eta=O(\frac{D}{\sqrt{M}})$ to minimize the upper bound to get the tightest upper bound:
$\sum_{m=1}^{M}\mathbb{E}[\langle\nabla F(z_{(k-1)M+m}),\Delta_{(k-1)M+m}-u_{k}\rangle]\leq O(D\sqrt{M}+MD\alpha)$ (75)
Part 2: bounding the regret of online gradient descent in eq. 49
For the second term in (49), we choose $u_{k}$ strategically to extract the Goldstein subdifferential:
$u_{k}=-D\cdot\frac{\sum_{m=1}^{M}\nabla F(z_{(k-1)M+m})}{\|\sum_{m=1}^{M}\nabla F(z_{(k-1)M+m})\|}$ (76)
With this choice of $u_{k}$:
$\sum_{m=1}^{M}\langle\nabla F(z_{(k-1)M+m}),u_{k}\rangle$ $=-D\cdot\left\|\sum_{m=1}^{M}\nabla F(z_{(k-1)M+m})\right\|$ (77)
$=-DM\cdot\left\|\frac{1}{M}\sum_{m=1}^{M}\nabla F(z_{(k-1)M+m})\right\|$ (78)
Substituting (75) and (78) into (49), and then into (48):
$F(x_{0})-\inf F$ $\geq\sum_{k=1}^{K}\left[-O(D\sqrt{M})-O(MD\alpha)+DM\cdot\left\|\frac{1}{M}\sum_{m=1}^{M}\nabla F(z_{(k-1)M+m})\right\|\right]$ (79)
Solving for the average over $k$:
$\frac{1}{K}\sum_{k=1}^{K}\left\|\frac{1}{M}\sum_{m=1}^{M}\nabla F(z_{(k-1)M+m})\right\|$ $\leq\frac{F(x_{0})-\inf F}{DMK}+O(\frac{1}{\sqrt{M}})+O(\alpha)$ (80)
For the randomly chosen output $x_{\text{out}}\sim\text{Uniform}\{x_{1},\ldots,x_{K}\}$:
$\mathbb{E}\left[\left\|\frac{1}{M}\sum_{m=1}^{M}\nabla F(z_{(k-1)M+m})\right\|\right]$ $\leq\frac{F(x_{0})-\inf F}{DMK}+O(\frac{1}{\sqrt{M}})+O(\alpha)$ (81)
The key insight is that these averages approximate the Goldstein subdifferential. Since $\|z_{(k-1)M+m}-x_{k}\|\leq MD\leq\delta$ (by our choice of $M=\lfloor\frac{\delta}{D}\rfloor$), we have:
$\nabla F(z_{(k-1)M+m})\in\partial_{\delta}F(x_{k})\text{ for all }m\in[M]$ (82)
By convexity of the Goldstein subdifferential:
$\frac{1}{M}\sum_{m=1}^{M}\nabla F(z_{(k-1)M+m})\in\partial_{\delta}F(x_{k})$ (83)
Therefore, from (81) and (83):
$\mathbb{E}[\text{dist}(0,\partial_{\delta}F(x_{\text{out}}))]$ $\leq\frac{F(x_{0})-\inf F}{DMK}+\Theta\left(\frac{1}{\sqrt{M}}\right)+\Theta(\alpha)$ (84)
To achieve $\mathbb{E}[\text{dist}(0,\partial_{\delta}F(x_{\text{out}}))]=\Theta(\epsilon)$, we set $\alpha=\Theta(\epsilon)$ and balance the remaining terms:
$\frac{F(x_{0})-\inf F}{DMK}+\Theta\left(\frac{1}{\sqrt{M}}\right)\leq\epsilon$ (85)
Let $C_{0}=F(x_{0})-\inf F$. We need both terms to be $\Theta(\epsilon)$:
$\frac{C_{0}}{DMK}$ $=\Theta(\epsilon)$ (86)
$\frac{1}{\sqrt{M}}$ $=\Theta(\epsilon)$ (87)
From (87), we get:
$\frac{1}{\sqrt{M}}=\Theta(\epsilon)\implies M=\Theta\left(\frac{1}{\epsilon^{2}}\right)$ (88)
Since $M=\lfloor\frac{\delta}{D}\rfloor$, we have $M\approx\frac{\delta}{D}$, which gives us:
$\frac{\delta}{D}=\Theta\left(\frac{1}{\epsilon^{2}}\right)\implies D=\Theta\left(\delta\epsilon^{2}\right)$ (89)
Let’s set $D=\Theta(\delta\epsilon^{2})$ and $M=\Theta\left(\frac{1}{\epsilon^{2}}\right)$ to satisfy this constraint. From (86), we can determine $K$:
$\frac{C_{0}}{DMK}=\Theta(\epsilon)\implies K=\Theta\left(\frac{C_{0}}{DM\epsilon}\right)$ (90)
Substituting our choices for $D$ and $M$:
$K$ $=\Theta\left(\frac{C_{0}}{\delta\epsilon^{2}\cdot\frac{1}{\epsilon^{2}}\cdot\epsilon}\right)$ (91)
$=\Theta\left(\frac{C_{0}}{\delta\epsilon}\right)$ (92)
Let’s set $K=\Theta\left(\frac{C_{0}}{\delta\epsilon}\right)$ to satisfy this constraint. For the step size $\eta$, we need to ensure stability of the algorithm. Based on standard analysis of stochastic gradient methods, we typically set:
$\eta=\Theta\left(\frac{D}{\sqrt{M}}\right)=\Theta\left(\delta\epsilon^{2}\cdot\epsilon\right)=\Theta\left(\delta\epsilon^{3}\right)$ (93)
Therefore, our final parameter settings are:
$D$ $=\Theta(\delta\epsilon^{2})$ (94)
$M$ $=\Theta\left(\frac{1}{\epsilon^{2}}\right)$ (95)
$K$ $=\Theta\left(\frac{C_{0}}{\delta\epsilon}\right)$ (96)
$\eta$ $=\Theta(\delta\epsilon^{3})$ (97)
Therefore, these parameter choices lead to $\mathbb{E}[\text{dist}(0,\partial_{\delta}F(x_{\text{out}}))]\leq\epsilon+O(\alpha)$